# Analyzing Games in Maker Protocol Part One: A Multi-Agent Influence Diagram Approach Towards Coordination

### v1

@@ 3.1 Mathematical Exposition
Definition 2
A node $D^{\prime}$ is $s$-reachable from a node $D$ in a MAID $K$ if there is some utility node $U_{e}$ such that if a new parent $D^{\prime}$ were added to $D$, there would be an active path in $K$ from $D$ to $U_{e}$ given $\phi(D)\subseteq j\lor\forall r$, where a path is active in a MAID if it is active in the same graph, viewed as a Bayesian Network.
Definition 3
The relevance graph for a MAID $K$ is a directed graph whose nodes are the decision nodes of $K$, and which contains an edge ${}_{\to}D^{\prime}$ if and only if $D^{\prime}$ is $s$-reachable from $D$.
Definition 4
If $K$ is a MAID and $C$ is a strategy profile for $K$, then the joint distribution for $K$ induced by $C$, denoted $Q_{K}^{C}$, is the joint distribution over $\mathcal{X}$ defined by the Bayes network where:
  • The set of variables is $\mathcal{X}$.
  • For each chance variable $X_{i}\in\mathcal{X}$, for each instantiation $\mathbf{pa}$ of parents of $X_{i}$, denoted $Pa(X_{i})$, there is an edge $X_{i}\rightarrow X_{i}$ iff $X_{i}\in\mathbf{pa}$.
  • For each decision variable $D_{i}\in\mathcal{D}$, the CPD for $D_{i}$ is $P(D_{i}|\mathbf{pa})$.
  • For each agent $A_{i}$, the CPD for each decision variable $D_{i}$ is specified by the corresponding decision rule in $C$.
We can now formulate an equation representing the expected utility that agent $A_{i}$ anticipates in a MAID $K$ when the agents adhere to a specific strategy profile $C$. Let $U_{A_{i}}^{C}$ denote the utility, and suppose $U_{A_{i}}^{C}=\sum_{j}u_{A_{i}}^{j}$. Then:
$E[U_{A_{i}}^{C}]=\sum_{\mathbf{v}\in\text{dom}(\mathbf{V})}u_{A_{i}}(\mathbf{v})\prod_{D\in\mathcal{D}}P(D|\mathbf{pa})\quad\text{where}\quad\text{dom}(\mathbf{V})=\text{joint domain of}\ \mathbf{V}.$ (1)
Since the expectation of a sum of random variables equals the sum of the expectations of the individual random variables, we can express this equation as:
$E[U_{A_{i}}^{C}]=\sum_{\mathbf{v}\in\text{dom}(\mathbf{V})}\sum_{j}u_{A_{i}}^{j}P(D|\mathbf{pa})\quad\text{where}\quad\text{dom}(\mathbf{V})=\text{joint domain of}\ \mathbf{V}.$ (2)
After defining the expected utility, we can now elucidate what it signifies for an agent to optimize their decision at one or more of their decision nodes, relative to a given set of decision rules for the other variables.
Definition 5
Let $H$ be a subset of $\mathcal{D}$, and let $C$ be a strategy profile. We define $C_{\text{opt}}$ as optimal for the strategy profile $C$ if, in the induced MAID $K_{(LCoI_{\text{opt}}M)}$, where the only remaining decisions are those in $H$, the strategy $C_{\text{opt}}$ is optimal. Formally, for all strategies $C_{\text{alt}}$:
$EU_{A_{i}}^{C_{\text{opt}}}$ $\geq EU_{A_{i}}^{C_{\text{alt}}}$
$EU_{A_{i}}^{C}$ $\geq EU_{A_{i}}^{C_{\text{opt}}}$
Note that, in this definition, the decision rules assigned to the variables in $H$ by $C$ are not relevant.
In the context of game theory, a strategy profile is deemed to represent rational behavior if it constitutes a Nash equilibrium [7]. Essentially, a strategy profile is a Nash equilibrium if no agent has an incentive to deviate from the strategy prescribed for them by the profile, provided that the other agents adhere to their specified strategies.
Definition 6
A strategy profile $C$ is a Nash equilibrium for a MAID $K$ if, for all agents $A_{i}$, $C_{\text{opt}}$ is optimal for the strategy profile.
3.2 Model Set Up
We analyse the coordination game played between the key stakeholders within the Maker Protocol : the governance DAO and users of the protocol. The dynamics of the game are as follows :
  1. 1.
    Governance entity (DAO), represented as a set of agents $\textbf{G}=\{G_{1},G_{2}\cdots G_{i}\}$, make decisions regarding protocol parameters, such as interest rates, collateral types, and governance policies.
  2. 2.
    Users of the protocol, represented as another set of agent $\textbf{Up}=\{Up_{1},Up_{2}\cdots Up_{i}\}$, make decisions regarding paying back or borrowing more Dai by adding collateral.
Referring [4], in the game between governance and users within the Maker Protocol, users interact with the protocol by borrowing Dai against collateral and governance entities, including voters and decision-makers, shape the protocol’s direction by proposing and voting on changes to parameters and policies. Users are incentivized to maintain adequate collateralization levels and repay debt obligations while governance entities seek to make decisions that enhance protocol stability, security, and efficiency while considering the preferences and welfare of the broader community.
3.3 Multi Agent Influence Diagram Construction
As per [13], the key risk parameters for Maker Vaults are:
  • Debt Ceiling: The Debt Ceiling represents the upper limit on the aggregate debt that can be generated through a singular collateral type within the Maker Protocol. Maker Governance allocates a specific Debt Ceiling to each collateral type, thereby facilitating adequate diversification within the collateral portfolio of the Maker Protocol. Upon reaching its designated Debt Ceiling, further creation of debt becomes unattainable unless existing users repay some or all of their outstanding Vault debt.
  • Stability Fee: The Stability Fee constitutes an annualized interest rate (APY) assessed on the total amount of Dai generated against a Vault’s collateral. This fee is exclusively denominated in Dai.
  • Liquidation Ratio: The Liquidation Ratio signifies the minimum collateral-to-debt ratio required within the Maker Protocol. A lower Liquidation Ratio implies an anticipation of relatively stable collateral prices, while a higher ratio indicates an expectation of greater price volatility. Mathematically, the Liquidation Ratio can be expressed as the ratio of the value of collateral to the total outstanding debt.
  • Liquidation Penalty: The Liquidation Penalty represents an additional fee imposed on the total outstanding Dai generated by a Vault upon liquidation. This penalty serves as an incentive for Vault owners to maintain adequate levels of collateral to mitigate the risk of liquidation. It is typically expressed as a percentage of the total debt outstanding.
  • Collateral Auction Duration: The Collateral Auction Duration specifies the maximum period allocated for the auctioning of collateral associated with Maker Vaults. While debt and surplus auction durations are system-wide parameters, the collateral auction duration is specific to individual Vaults.
  • Auction Bid Duration: The Auction Bid Duration denotes the duration for which an individual bid remains valid before expiring and subsequently closing the auction. This parameter is crucial in determining the efficiency and speed of the auction process within the Maker Protocol.
  • Auction Step Size: The Auction Step Size serves as a risk parameter aimed at incentivizing early participation in auctions and preventing potential abuse by discouraging minimal bid increments. It establishes the minimum bid increment required for successive bids in an auction, promoting a fair and efficient auction mechanism.
3.3.1 Smart contract Modules:
@@
The Maker Protocol relies on various smart contracts to automate its core functionalities. These modules interact with each other and with external actors to maintain system stability and facilitate user interactions. As per [16] :
@@
The below figure shows the MAID for this game :
Multi Agent Influence Diagram for GovernanceDAO and User game
Deeper Influence Diagram to show interdependencies between parameters on the smart contract level
The MAID model captures the dependencies and influences between the decisions of the governance entity and the users of the Maker Protocol. Nodes in the MAID represent variables such as decisions, uncertainties, utilities, and influences, while directed edges indicate causal relationships or dependencies between them.
The game unfolds over discrete time steps, with each agent making sequential decisions based on their objectives, beliefs, and the actions of the other agent. The decisions of one agent influence the payoffs and strategies available to the other agent, creating a dynamic interplay of incentives and behaviors.
A brief overview of the connections is as follows :
Sustainability
The sustainability of the Maker Protocol hinges on prudent management of risk parameters such as the debt ceiling, liquidation ratio, and stability fee. While higher debt ceilings can potentially increase revenue through stability fees, they also expose the protocol to greater risk if not managed carefully. Similarly, a lower liquidation ratio allows for more borrowing and higher fee generation but escalates the risk of undercollateralization and losses during market downturns. The stability fee, crucial for revenue generation, must strike a delicate balance to avoid discouraging borrowing and reducing demand for DAI. The Governance Module assumes a pivotal role in aligning decisions regarding risk parameters and protocol upgrades with sustainability goals.
Profitability
Profitability within the Maker Protocol ecosystem is intricately linked to the effectiveness of risk parameters and smart contract modules. The stability fee, determined by governance, directly impacts the profitability of Vault owners by influencing the cost of borrowing Dai against collateral. Additionally, the efficacy of the System Stabilizer Module in managing liquidations and surplus auctions affects the profitability of participants, including Keepers engaging in these auctions to earn profits. Higher liquidation penalties incentivize users to maintain adequate collateralization, thereby reducing the need for liquidations and potential losses for the protocol.
Risk Parameters
Risk parameters, governed by the Governance Module, play a pivotal role in capital efficiency, risk mitigation, and profitability within the Maker Protocol ecosystem. Parameters such as the liquidation ratio, debt ceiling, liquidity penalty, and stability fee directly influence user debt, security risk, demand for DAI, and collateral dynamics. These parameters must be meticulously adjusted to ensure high capital efficiency, low risk exposure, and sustainable profitability. For instance, the debt ceiling sets limits on debt generation against specific collateral types to prevent over-leveraging and maintain system stability. Similarly, the liquidation ratio and penalty incentivize users to uphold sufficient collateralization levels, mitigating the risk of liquidation events.
Debt
The debt within the Maker Protocol represents borrowed Dai against provided collateral. Smart contract modules like the Core Module and Rate Module are instrumental in managing user debt by maintaining accurate records and dynamically adjusting stability fees based on market conditions. Risk parameters such as the debt ceiling, liquidation ratio, and penalty further influence debt dynamics by setting limits and incentivizing adequate collateralization.
Demand for DAI
The demand for DAI is intricately linked to borrowing conditions within the Maker Protocol, influenced primarily by the stability fee. Lower stability fees make borrowing DAI more attractive, potentially increasing demand, although this must be balanced with the need for protocol revenue. Additionally, risk parameters such as the liquidation ratio and debt ceiling play crucial roles in managing risk and ensuring the stability of DAI, thus impacting demand. The efficiency of collateral auctions facilitated by modules like the System Stabilizer Module also influences confidence in the protocol and, consequently, demand for DAI.
Security Risks
Security risks within the Maker Protocol stem from vulnerabilities in smart contracts, market manipulation, or unexpected events like oracle failures. The Oracle Module is vital for mitigating security risks by providing accurate price feeds, while the Maker Protocol Emergency Shutdown mechanism serves as a safeguard against critical failures. Adjusting parameters like the liquidation ratio and debt ceiling is essential for managing security risks associated with insolvency or exposure to volatile assets.
Market Conditions
@@
Market conditions significantly impact the stability and performance of the Maker Protocol. Risk parameters such as the liquidation ratio and penalty are designed to account for market fluctuations and maintain system health during adverse conditions. The Oracle Module provides real-time price data, enabling swift reactions to changing market conditions, while dynamically adjusting debt ceilings based on market conditions helps manage risk and maintain stability.
Collateral Dynamics
@@
The Collateral Module manages collateral types and associated parameters within the Maker Protocol. Adjusting risk parameters like the debt ceiling and liquidation ratio influences collateral dynamics by affecting the selection and utilization of collateral assets. Furthermore, the efficiency of collateral auctions, facilitated by modules like the System Stabilizer Module, impacts collateral dynamics by determining the ease of collateral liquidation in case of undercollateralization.
4 Conclusion and Future Work
In this paper, we have successfully quantified the different inter dependencies between the different parameters in the Maker protocol and created a MAID like structure to analyse the games in the protocol. In future iterations, we will analyse the game using the mathematical principles laid down by Koller and Milch and provide a comprehensive view of equilibria in the protocol in different conditions. In subsequent papers, we will also show that MAIDs can be used as a way to assess the incentive security of a defi protocol such as Maker. Using the library developed by Google Deepmind (see [14]), we will develop our own code with a thorough analysis of collateral dynamics (see [15] ), quantify expected values and payoff functions using ideas from Section 3, and we will demonstrate the equilbrium conditions of the protocol in real time so as to motivate a formal verification procedure of game theoretic interactions on Maker protocol.

### v2

@@ 3. Model
We define utilities as follows: For agent $a\in\mathcal{A}$,
(3) $u_{a}(d_{a}|\texttt{CR}=\texttt{risky},\texttt{MS}=\texttt{bad})$ =$=\begin{cases}+100&\text{ if }d_{a}=\texttt{yes},\\ -100&\text{ if }d_{a}=\texttt{no}.\end{cases}$
(4) $u_{a}(d_{a}|\texttt{CR}=\texttt{risky},\texttt{MS}=\texttt{good})$ =$=\begin{cases}+50&\text{ if }d_{a}=\texttt{yes},\\ -50&\text{ if }d_{a}=\texttt{no}.\end{cases}$
(5) $u_{a}(d_{a}|\texttt{CR}=\texttt{not risky},\texttt{MS}=\texttt{bad})$ =$=\begin{cases}+25&\text{ if }d_{a}=\texttt{yes},\\ -25&\text{ if }d_{a}=\texttt{no}.\end{cases}$
(6) $u_{a}(d_{a}|\texttt{CR}=\texttt{not risky},\texttt{MS}=\texttt{good})$ =$=\begin{cases}-100&\text{ if }d_{a}=\texttt{yes},\\ +100&\text{ if }d_{a}=\texttt{no}.\end{cases}$
3.1. Explanation of the Utilities
As we have defined two chance variables—collateral risk and market sentiment—we need distinguish between four cases.
  1. (1)
    For any agent $a\in\mathcal{A}$, given a collateral asset classified as risky and a prevailing negative market sentiment, voting in favor of the proposal (decision variable $d_{a}=\texttt{yes}$) yields a positive of $+100$. Conversely, voting against the proposal ($d_{a}=\texttt{no}$) results in a utility of $-100$.
  2. (2)
    For any agent $a\in\mathcal{A}$, given a risky collateral asset and a positive market sentiment, voting in favor of the proposal (decision variable $d_{a}=\texttt{yes}$) results in a positive utility of $+50$. Conversely, voting against the proposal ($d_{a}=\texttt{no}$) leads to a utility of $-50$.
  3. (3)
    For any agent $a\in\mathcal{A}$, given a non-risky collateral asset and a negative market sentiment, voting in favor of the proposal (decision variable $d_{a}=\texttt{yes}$) yields a positive utility of $+25$, primarily driven by anticipated future risk mitigation benefits. Analogously for $d_{a}=\texttt{no}$.
  4. (4)
    For any agent $a\in\mathcal{A}$, given a non-risky collateral asset and a positive market sentiment, voting in favor of the proposal (decision variable $d_{a}=\texttt{yes}$) is deemed unsustainable and results in a utility decrement of 100. Analogously for $d_{a}=\texttt{no}$.
Remark 2.
We consider equal voting power regardless of the amount of \$GOV tokens that each agent has. Under stake-based voting, each agent’s expected utility would be scaled by the proportion of \$GOV tokens. In our model, we assume independence between the chance variables—market sentiment and collateral risk. We do not assume a direct causal link between market sentiment and collateral risk (Pearl, 2009). We assume homogeneous agents to make the model tractable.
4. Analysis
As devised in Section 2, we present some experimental results using the ”pycid” library developed by Google Deepmind (Fox et al., 2021). We start by initialising and devising dummy variables as follows:
[…]
  • Recall market sentiment as MS. For $\texttt{MS}\in\{\texttt{good},\texttt{bad}\}$, we similarly create dummy binary variables where good is denoted by 0 and bad is denoted by 1. Again, when we say that market sentiment is ”good”, we mean that the market conditions for the collateral asset and future sentiment for its price volatility is favourable for long term investment and vice versa when we use the term ”bad”.
  • Vote as V. Notice that $\texttt{V}=d_{a}$ and the notation was chosen for the slight difference to represent the action of voting and for ease of keeping track. Recall that $\texttt{V}=d_{a}=\{\texttt{yes},\texttt{no}\}=\{\texttt{1},\texttt{0}\}$ for $a\in\mathcal{A}$. Simply, yes implies voting for the proposal and no is voting against the proposal. Every agent in the protocol has the same exact choices for the proposal.
  • Utility as U as defined above as a utility function.
4.1. Single Agent Optimal Strategy
We have all the tools that we need to analyse the model. Let us consider the case of a single agent in this model of governance. Given inputs$(\mathcal{A},\chi,\mathcal{D},\mathcal{U})$ as we have defined above in Section 3 equations $(1),(2),(3)$ and $(4)$, we have all the ingredients required to develop the MAID for the single agent in question and the dynamics of the governance protocol as defined above can be summarized in the causal influence diagram (CID) in Figure 1.
Figure 1. Causal Influence Diagram for Model
[…]
Under the utility functions defined in the previous section and using the preliminaries, the optimal policy for $V=0$ (Vote against) and $V=1$ (Vote For) can be summarised in Figure 2.
Figure 2. Optimal Policy for an Agent in the Game
Using definition of optimal strategy, for any agent $a\in\mathcal{A}$,
$EU_{a}(\sigma_{V}(1))\geq EU_{a}(\sigma_{V}(0))$
whenever $\texttt{CR}=1$ or $\texttt{MS}=1$. Similarly
$EU_{a}(\sigma_{V}(0))\geq EU_{a}(\sigma_{V}(1))$
when $\texttt{CR}=0$ and $\texttt{MS}=0$ That is, for any arbitrary agent in the MAID, the optimal strategy $\sigma$ for every agent would be to vote for strengthening the risk parameters $(\texttt{V}=1)$ in case the collateral dynamics are risky $(\texttt{CR}=1)$ or if the market sentiment is bad $(\texttt{MS}=1)$and vote against the proposal $(\texttt{V}=0)$ only if both the collateral dynamics are less volatile $(\texttt{CR}=0)$ and the market sentiment is good $(\texttt{MS}=0)$ too. Figure 2 gives a good summary of this as an output using the pycid library. The expected utility for every single arbitrary agent is 68.75, solved as per the preliminaries and the specifics of the model i.e. $EU(\sigma)=\sum_{U\in\mathcal{U}_{a}}\sum_{u\in dom(U)}P_{\mathcal{M}[\sigma]}(U=u)\cdot u={0.25}\cdot({100+50+25+100}){=68.75}$. Notice that in case the voting power was stake based, the expected utility would have been \$GOV owned by agent\$GOV total $\frac{\texttt{\$GOV owned by agent}}{\texttt{\$GOV total supply}}\cdot\texttt{68.75}$
4.2. Multi Agent Equilibrium
To put things back into perspective, we know the optimal strategy for an agent in our governance model and so we have an understanding of their independent decisions when faced with the same conditions. But what does it mean for their collective equilibrium? Let us formally address this.
Remark 3.
If we assume that all agents in the governance protocol are making the decisions simultaneously, then it is safe to also assume that their independent decisions would converge to the same decision. So we make another assumption of non simultaneous decisions and every agent can see each other’s vote to add in some complexity.
Remark 4.
In a DeFi governance model where agents are homogeneous and share the same optimal strategy but make decisions at different times, voting on a proposal takes the form of a Stackelberg Game, where the leader acts according to their optimal policy first.
To see this in action let us consider the case of two agents in the system and their interactions. Since the population of governance voters is assumed to be homogeneous and similar, taking case of two agents is sufficient. Like before let us define and initialise variables as follows:
  • Agent 1 as A1
  • Agent 2 as A2
  • Utility of Agent 1 as U1
  • Utility of Agent 2 as U2
We change the perspective of the game. As we have assumed before, every agent $a\in\mathcal{A}$ has the same CID $(\mathcal{V}_{a},E_{a})$ which implies the same chance $\chi$, decision variables $\mathcal{D}_{a}$ and utility variables $\mathcal{U}_{a}$ for all agents.
Given some value of chance variables CR and MS, agent A1 has made a decision to vote V on the proposal to derive utility U1 which will be seen by the other agent in the governance protocol. The influence diagram (Figure 3) implies that each agent’s decision has an effect on the other’s utility. Recall the non simultaneous criterion that implies that one agent’s vote influences the other agent’s vote and hence also the decision.
Figure 3. Multi Agent Causal Influence Diagram
[…]
In all other cases we claim that the utilities would be $\leq 0$. Indeed, since agents A1 and A2 are following their optimal strategies and are homogeneous (similar), a positive utility U2 for A2 by choosing a different decision than A1 breaks the homogeneous assumption since that would imply that one of the two agents derive more utility from a decision than the other which is not possible.
Remark 5.
Given a system with agents A1 and A2, where A1 is designated as the leader, if agent A2 experiences a net positive utility ($U_{2}>0$) by adopting a strategy divergent from A1’s, the following conditions must hold:
  1. (1)
    Agent Heterogeneity: The utility functions of A1 and A2 are not identical, implying diverse preferences or objectives.
  2. (2)
    Strategic Rationality: Agent A2’s decision to deviate is a strategic choice, prioritizing its individual utility maximization over collective welfare or protocol adherence.
Under these conditions, A2’s behavior can be characterized as strategic or potentially adversarial, posing a risk to the system’s equilibrium.
Using Algorithm 1, we obtain the results for the pure Nash equilibrium of the system displayed in Figure 4.
Figure 4. Pure Nash Equilibrium for Two Agents
The analysis indicates that in a two-agent system where agent A1 acts as a leader, the optimal strategy for agent A2 is to mimic A1’s decision. This outcome aligns with the equilibrium of a Stackelberg game (Julien, 2018), wherein the leader’s choice precedes and influences the follower’s action. Consequently, within a homogeneous agent population, where all agents possess identical utility functions and information, there exists a strong incentive for each agent to adopt the optimal policy of every other agent, leading to a collective action scenario. This observation supports the proposed hypothesis.
What does it mean for DeFi governance protocols? One could claim that if the optimal strategy for any arbitrary agent in a DeFi governance is to promote protocol sustainability then the protocol should have no instances of manipulation (such as Governance Extractable Value (Werner et al., 2022)) assuming homogeneity of agents. In case there is a deviation from the optimal strategy (or if there exists an incentive to deviate) then there will exist an instance of ”attack” behavior which may be harmful to the protocol. In the next section, we look at a very simple attack model using MAIDs.
4.3. A Simple Attack Model
Previously, our model assumed homogeneous agents. We now introduce an adversarial agent to capture potential malicious behavior. Formally, we extend the model in Section 3 by introducing an additional agent $a^{\prime}\in\mathcal{A}$ with identical chance variables $\chi$ and decision space $\mathcal{D}_{a^{\prime}}=\mathcal{D}_{a}$ as the other agents. However, $a^{\prime}$ possesses a distinct utility function $u_{a^{\prime}}$. This adversarial agent aims to disrupt the protocol by acting contrary to the interests of honest agents.
Thus the utility function of agent $a^{\prime}$ can be written out as follows:
(7) $u_{a^{\prime}}(d_{a^{\prime}}|\texttt{CR}=\texttt{risky},\texttt{MS}=\texttt{bad})$ =$=\begin{cases}-100&\text{ if }d_{a^{\prime}}=\texttt{yes},\\ +100&\text{ if }d_{a^{\prime}}=\texttt{no}.\end{cases}$
(8) $u_{a^{\prime}}(d_{a^{\prime}}|\texttt{CR}=\texttt{risky},\texttt{MS}=\texttt{good})$ =$=\begin{cases}-50&\text{ if }d_{a^{\prime}}=\texttt{yes},\\ +50&\text{ if }d_{a^{\prime}}=\texttt{no}.\end{cases}$
(9) $u_{a^{\prime}}(d_{a^{\prime}}|\texttt{CR}=\texttt{not risky},\texttt{MS}=\texttt{bad})$ =$=\begin{cases}-25&\text{ if }d_{a^{\prime}}=\texttt{yes},\\ +25&\text{ if }d_{a^{\prime}}=\texttt{no}.\end{cases}$
(10) $u_{a^{\prime}}(d_{a^{\prime}}|\texttt{CR}=\texttt{not risky},\texttt{MS}=\texttt{good})$ =$=\begin{cases}+100&\text{ if }d_{a^{\prime}}=\texttt{yes},\\ -100&\text{ if }d_{a^{\prime}}=\texttt{no}.\end{cases}$
Consider a governance proposal to strengthen risk parameters for a risky collateral asset. Under standard assumptions of agent homogeneity, the optimal strategy and Nash equilibrium for agents is to vote ”yes” ($d_{a}=\texttt{yes}$), yielding a utility of $+100$.
We introduce an adversarial agent $a^{\prime}\in\mathcal{A}$ with the primary objective of disrupting the protocol. In contrast to honest agents, $a^{\prime}$ optimally votes ”no” ($d_{a^{\prime}}=\texttt{no}$), resulting in a utility of $+100$.
Formally, the system comprises two agents, $a$ and $a^{\prime}$, representing honest and adversarial actors, respectively. Assuming identical CID and MAID structures as in Section 4.2, the system’s architecture aligns with Figure 3 where $A_{2}=a^{\prime}$ and $A_{1}=a$.
where U1 and U2 are the utilities of the honest agent $a$ and attacker $a^{\prime}$ respectively. We construct a payoff matrix for both the agents through their utilities given a decision that they make about the proposal given above. The payoff matrix can be summarised as follows:
a/a’ 0 1
0 0,100 -100,-100
1 100,100 100,-100
The Nash equilibrium for this set-up using Algorithm 1 is given in Figure 5.
Figure 5. Nash Equilibrium for Honest Agent vs Attacker
Solving the MAID reveals that agent $a^{\prime}$ adopts a dominant strategy of choosing 0, irrespective of the decision of $a$. Conversely, $a$ exhibits a dominant strategy of selecting 1, independent of the choice of $a^{\prime}$. Both agents have an expected utility of 0. Honest agents achieve non-negative net utility when constituting a majority, whereas adversarial agents attain non-negative utility under majority control. The application of MAIDs effectively demonstrates the framework’s capacity to formally verify agent incentives within a governance protocol, highlighting its accessibility and utility for analyzing strategic interactions.
5. Application to MakerDAO Governance Protocol
The Maker Protocol operates as a decentralized finance (DeFi) platform on the Ethereum blockchain, facilitating the creation of Dai, a stablecoin pegged to the US dollar. As per (MakerDAO, 2024), Users engage with the protocol by depositing cryptocurrency assets, such as Ethereum or other tokens, into smart contracts as collateral. Based on the value of this collateral, users can then generate Dai, typically up to a predetermined percentage of the collateral’s worth. To maintain the stability and security of the system, the protocol enforces various risk parameters, including collateralization ratios and debt ceilings, which dictate the relationship between collateral and the amount of Dai that can be generated. Users pay stability fees on the Dai they generate, serving as interest rates that contribute to the protocol’s stability. Governance of the Maker Protocol is decentralized, with MKR token holders participating in decision-making through a voting mechanism. These holders vote on proposals to adjust risk parameters and other protocol settings, ensuring the protocol remains responsive to market conditions and emerging risks. Additionally, the protocol includes mechanisms for liquidating collateral if its value falls below a certain threshold, helping to mitigate systemic risk and maintain the stability of the Dai ecosystem. More information about the governance dynamics of the Maker protocol can be found in (Kiayias and Lazos, 2022).
The binary voting mechanism employed by MakerDAO, where MKR holders cast yes/no votes on proposals, is analogous to the governance games analyzed in preceding sections. By applying MAIDs to a simplified binary voting model, we can examine agent interactions under two scenarios: homogeneous agents sharing identical MAIDs and heterogeneous agents with opposing objectives (e.g., protocol sustainability vs. disruption). This analysis provides insights into the game-theoretic dynamics within the protocol and MAIDs can be used in similar ways to model more complex interactions and in turn be used to create a formal verification system for coordination games in DeFi governance protocols.
5.1. MAID for MakerDAO
In this section, we outline the structure of a game between the users of the Maker protocol and the MakerDAO governance member on the smart contract level. For the MAID, the variables are as follows:
  • The chance variables $\chi$ include smart contract modules, market conditions, collateral dynamics, demand for DAI stablecoin and security risks.
  • The decision variables $\mathcal{D}$ include debt (that the user wants to undertake) and the risk parameters (that the governance DAO members want to set for a particular asset).
  • The utility variables $\mathcal{U}$ include profitability for the users of the protocol and protocol sustainability for the governance members of the DAO.
The MAID is depicted in Figure 6.
Figure 6. MAID for Maker on the Smart Contract Level
5.1.1. Chance Variables
@@
First, we outline the various smart contracts modules (Foundation, 2014) that interact with each other and with external actors to maintain system stability and facilitate user interactions. Then, we describe the remaining chance variables.
  1. (1)
    Smart Contract Modules:
@@
  2. (2)
@@
    Market Conditions: Market conditions significantly impact the stability and performance of the Maker Protocol. Risk parameters such as the liquidation ratio and penalty are designed to account for market fluctuations and maintain system health during adverse conditions. The Oracle Module provides real-time price data, enabling swift reactions to changing market conditions, while dynamically adjusting debt ceilings based on market conditions helps manage risk and maintain stability.
  3. (3)
@@
    Collateral Dynamics: The Collateral Module manages collateral types and associated parameters within the Maker Protocol. Adjusting risk parameters like the debt ceiling and liquidation ratio influences collateral dynamics by affecting the selection and utilization of collateral assets. Furthermore, the efficiency of collateral auctions, facilitated by modules like the System Stabilizer Module, impacts collateral dynamics by determining the ease of collateral liquidation in case of undercollateralization.
  4. (4)
    Demand for DAI: The demand for DAI is intricately linked to borrowing conditions within the Maker Protocol, influenced primarily by the stability fee. Lower stability fees make borrowing DAI more attractive, potentially increasing demand, although this must be balanced with the need for protocol revenue. Additionally, risk parameters such as the liquidation ratio and debt ceiling play crucial roles in managing risk and ensuring the stability of DAI, thus impacting demand. The efficiency of collateral auctions facilitated by modules like the System Stabilizer Module also influences confidence in the protocol and, consequently, demand for DAI.
  5. (5)
    Security Risks: Security risks within the Maker Protocol stem from vulnerabilities in smart contracts, market manipulation, or unexpected events like oracle failures. The Oracle Module is vital for mitigating security risks by providing accurate price feeds, while the Maker Protocol Emergency Shutdown mechanism serves as a safeguard against critical failures. Adjusting parameters like the liquidation ratio and debt ceiling is essential for managing security risks associated with insolvency or exposure to volatile assets.
5.1.2. Decision Variables
  1. (1)
    Debt: The debt within the Maker Protocol represents borrowed Dai against provided collateral. Smart contract modules like the Core Module and Rate Module are instrumental in managing user debt by maintaining accurate records and dynamically adjusting stability fees based on market conditions. Risk parameters such as the debt ceiling, liquidation ratio, and penalty further influence debt dynamics by setting limits and incentivizing adequate collateralization.
  2. (2)
    Risk Parameters: Risk parameters, governed by the Governance Module, play a pivotal role in capital efficiency, risk mitigation, and profitability within the Maker Protocol ecosystem. Parameters such as the liquidation ratio, debt ceiling, liquidity penalty, and stability fee directly influence user debt, security risk, demand for DAI, and collateral dynamics. These parameters must be meticulously adjusted to ensure high capital efficiency, low risk exposure, and sustainable profitability. For instance, the debt ceiling sets limits on debt generation against specific collateral types to prevent over-leveraging and maintain system stability. Similarly, the liquidation ratio and penalty incentivize users to uphold sufficient collateralization levels, mitigating the risk of liquidation events.
5.1.3. Utility Variables
  1. (1)
    Profitability: Profitability within the Maker Protocol ecosystem is intricately linked to the effectiveness of risk parameters and smart contract modules. The stability fee, determined by governance, directly impacts the profitability of Vault owners by influencing the cost of borrowing Dai against collateral. Additionally, the efficacy of the System Stabilizer Module in managing liquidations and surplus auctions affects the profitability of participants, including Keepers engaging in these auctions to earn profits. Higher liquidation penalties incentivize users to maintain adequate collateralization, thereby reducing the need for liquidations and potential losses for the protocol.
  2. (2)
    Sustainability: The sustainability of the Maker Protocol hinges on prudent management of risk parameters such as the debt ceiling, liquidation ratio, and stability fee. While higher debt ceilings can potentially increase revenue through stability fees, they also expose the protocol to greater risk if not managed carefully. Similarly, a lower liquidation ratio allows for more borrowing and higher fee generation but escalates the risk of undercollateralization and losses during market downturns. The stability fee, crucial for revenue generation, must strike a delicate balance to avoid discouraging borrowing and reducing demand for DAI. The Governance Module assumes a pivotal role in aligning decisions regarding risk parameters and protocol upgrades with sustainability goals.
6. Conclusion and Future Work
In this study, we explored a simple DeFi governance game and analyzed its equilibria in both single and multi-agent settings using Multi-Agent Influence Diagrams (MAIDs). Our analysis included a practical example involving an attacker and provided the framework for constructing a MAID specifically tailored to MakerDAO, which can serve as a foundation for further exploration of decentralized governance dynamics. Our work demonstrates the potential of MAIDs as a powerful tool for understanding the complex interactions within DeFi governance models. However, there are several promising directions for future research. Firstly, our current model employs discrete uniform conditional probability distributions; extending this to continuous distributions could yield more insights. Additionally, the analysis was limited to two agents, and there is significant potential to expand this framework to accommodate a larger number of agents, each with possibly different utility functions, to better mirror the diverse landscape of DeFi ecosystems. Moreover, while our study does not assume a direct causal link between market sentiment and collateral risk, incorporating such causation could provide a deeper understanding of the effects on governance outcomes. Solving the MAID for MakerDAO using real-world parameters remains an open challenge and will be crucial for validating the applicability of our model. . As DeFi continues to grow, such tools will be instrumental in guiding the development of secure and resilient governance structures.